# Quantized polarization in a generalized Rice-Mele model at arbitrary filling

### v1

@@ Abstract
We discuss the charge polarization in a generalized Rice-Mele model at arbitrary particle filling per site as a model of charge ordered systems in one dimension. The model possesses neither the conventional bond-centered inversion symmetry nor the one site translation symmetry alone, but has combinations of these symmetries. We show that the charge polarization in the ground state is quantized by the combined symmetry and is characterized solely by the filling. Especially, the polarization can be $1/2$ (mod 1) in the zero filling limit. Under the open boundary condition, there exist excess charges accumulated near edges of the system irrespective of existence or absence of edge modes. Correspondingly, we decompose the polarization into a bulk contribution and an edge contribution, and numerically demonstrate that the polarization is dominated by the former (latter) when the energy gap is large (small). We also discuss a simple generalization of our model and examine absence/existence of a gapless edge mode protected by the inversion symmetry by introducing intra unit cell and inter unit cell contributions of the charge polarization.
@@
An insulator can have polarization which characterizes asymmetry of a charge configuration. A classical example is the ionic insulator where positive point charges and negative point charges are alternatively distributed. Careful discussions are necessary for an appropriate definition of the polarization even for such a simple system under the periodic boundary condition. It is now widely accepted that the polarization can well be described by the Berry-Zak formula, which clarifies an intimate relationship of the polarization to topology and symmetry  Resta and Vanderbilt (2007); Vanderbilt (2018); Spaldin (2012); King-Smith and Vanderbilt (1993); Vanderbilt and King-Smith (1993); Zak (1989). The celebrated prototypical model of dipole topological insulators is the Su-Schrieffer-Heeger (SSH) model, where the hopping integral has a staggered modulation Su et al. (1979); Su and Schrieffer (1981); János K. Asbóth (2016). The polarization of the SSH model is quantized by an inversion symmetry, and there have been extensive theoretical studies of the generalized SSH models and many interesting properties have been clarified Li et al. (2014); Yao et al. (2017); Pérez-González et al. (2019); Scollon and Kennett (2020); Guo and Chen (2015); Martinez Alvarez and Coutinho-Filho (2019); Huda et al. (2020); Maffei et al. (2018); Xie et al. (2019). Topological nature of the charge polarization is clearly seen in charge pumping where a change of the polarization in an adiabatic cycle in a parameter space is quantized by the Chern number Thouless (1983); Citro and Aidelsburger (2023). Indeed, the Rice-Mele (RM) model Rice and Mele (1982) was experimentally realized in cold atoms and the Thouless pumping was observed Lohse et al. (2016); Nakajima et al. (2016, 2021). In this way, the SSH and RM models are regarded as one of the most important models in the context of topological insulators.
@@
The very basic physics behind the SSH and RM models is charge ordering due to the Peierls instability  Kagoshima (1981); van Smaalen (2004) as they were originally proposed for polyacetylene and diatomic polymers Su et al. (1979); Su and Schrieffer (1981); Rice and Mele (1982). A trimerized variant of the SSH model was also studied with an application to the Peierls system TTF-TCNQ Su and Schrieffer (1981), although related models have been discussed in relation to various systems such as engineered materials and cold atoms Guo and Chen (2015); Martinez Alvarez and Coutinho-Filho (2019); Huda et al. (2020); Maffei et al. (2018); Xie et al. (2019). In a general one-dimensional metal with the Fermi wavenumer $k_{F}$, the charge susceptibility shows singularity at the wavenumber $Q=2k_{F}$ and the system has strong instability toward a charge order. The Peierls phase transition and charge orders have been discussed in both organic materials and inorganic materials, and there are a variety of charge orders with different modulations $Q$  Kagoshima (1981); van Smaalen (2004). In this context, the standard SSH and RM models correspond to the simplest charge order with $Q=\pi$. In spite of the extensive studies on the charge polarization, however, focus has been made mainly on this particular case and polarizations for other charge orders with general $Q$-vectors have not been well explored. Since there are a lot of charge ordered one-dimensional materials Kagoshima (1981); van Smaalen (2004), it is natural to consider generalization of the $Q=\pi$ order to $Q=2k_{F}\neq\pi$ orders. Such generalization may be interesting also in the context of symmetry protected topological phases beyond the non-interacting fermions. For example, an $S=1/2$ trimer spin chain exhibits a 1/3 magnetization plateau implying a uniquely gapped ground state under a magnetic field, which can be regarded as a generalization of more familiar spin dimer orders Hida (1994); Hu et al. (2014); Bera et al. (2022). Modulated spin systems could have “charge polarization” associated with the U(1) spin rotation symmetry similarly to charge ordered fermion systems Oshikawa et al. (1997); Nakamura and Todo (2002); Tasaki (2018); Watanabe et al. (2021); Furuya and Sato (2021).
@@ II Quantized charge polarization under symmetry
where the sites are labeled as $j=0,1,2,\cdots,L-1$. The modulation wavenumber is given by $Q=2\pi\nu$, where $\nu=N/L=p/q$ is the particle filling per site with $N$ being the total number of fermions and $p,q$ being coprime integers ($p$ fermions per $q$-site unit cell on average). The system size is assumed to be an integer multiple of the unit cell for both the periodic and open boundary conditions, $L=L_{q}q$, and is taken to be $L=600$ for numerical calculations in this study. The modulation could be induced by interactions or arise from different atom species in real materials, but here $(\delta,\Delta)$ are model parameters of the system. Note that the modulation factor $\cos(Qj)$ reduces to the conventional factor $(-1)^{j}$ at the half-filling $\nu=1/2$. One can introduce a phase shift such as $\cos(Q(j-j_{0}))$ with $j_{0}=1,2,\cdots,q-1$ for which the corresponding Hamiltonian is related to the above $H_{0}$ by the $j_{0}$-site translation (see Sec. IV). It is straightforward to numerically calculate single-particle spectra of Eq. (1) and there exists an energy gap $\Delta E$ between the $N$-th and $(N+1)$-th single-particle energy eigenvalues under the periodic boundary condition as shown in Fig. 1. One can add interactions such as $H_{\rm int}=\sum_{ij}V_{ij}n_{i}n_{j}$ where $n_{j}=c_{j}^{\dagger}c_{j}$ and $V_{ij}$ keeps the symmetry of the system. A large part of our discussions including symmetry-based arguments presented below is applicable also to interacting systems, which will be important when one considers bosonic particles and spin systems.
@@
Let us discuss symmetry of the Hamiltonian (1) and the resulting quantization of the charge polarization under the periodic boundary condition. Obviously, neither the one site translation symmetry $T_{x}$ nor the bond-centered inversion symmetry $I_{0}$ about the bond $\braket{L-1,0}$ alone is preserved for general $Q$ when $\delta$ and $\Delta$ are non-zero. ($\braket{j,j+1}$ denotes the bond between the sites $j$ and $j+1$.) However, one can see that appropriate combinations of these symmetries are preserved. Indeed, the Hamiltonian is invariant under the combined symmetry when either $\delta$ or $\Delta$ is zero,
@@
The operator $U_{x}$ is often called Lieb-Schultz-Mattis twist operator  Oshikawa et al. (1997); Nakamura and Todo (2002); Bohm (1949); Lieb et al. (1961); Tasaki (2018, ) and its expectation value in a uniquely gapped ground state is known to be $|\braket{U_{x}}|\simeq 1$ in one dimension. This is because the energy of the variational state |$\ket{\Phi}=U_{x}\ket{\Psi}$ converges to the ground state energy in the thermodynamic limit and therefore |$\ket{\Phi}$ must be essentially proportional to the ground state with a phase factor, |$\ket{\Phi}\simeq e^{i2\pi P_{x}}\ket{\Psi}$, in a uniquely gapped system. See also Appendix A. Thus the Resta formula for the polarization is well-defined in one dimensional uniquely gapped ground state, although the expectation value is vanishing in the thermodynamic limit for general dimensions and an appropriate modification is necessary to resolve such a problem  Watanabe and Oshikawa (2018); Tada and Oshikawa (2023).
@@
This implies that, for the uniquely gapped ground state |$\ket{\Psi}$ which is an eigenstate of the combined operator $I_{k}$,
@@
In this section, we consider the model Eq. (1) under the open boundary condition where there is no hopping between the sites $j=L-1$ and $j=0$. In the standard SSH model ($\delta\neq 0,\Delta=0$) at $\nu=1/2$, there can be gapless edge modes thanks to the chiral symmetry, but existence of such single-particle edge modes are not generally guaranteed in a system only with point group symmetry Song et al. (2017); Huang et al. (2017). Similarly, there are no edge modes in the present system at almost all filling except for some special cases which will be discussed in Sec. IV. Nevertheless, one may naively expect that there are extra charges accumulated near edges of the system for a given filling. Here we introduce excess charges over one unit cell (length $q$) compared to the uniform background charges,
@@
Figure 5 shows examples of the fermion density profiles $\braket{n_{j}}$ and the corresponding excess charges $\delta n_{k}$. Clearly, the excess charges vanish in the bulk region of the system, although the fermion density itself oscillates with the period $q$. This is a general property for a gapped periodic system which has unit cell translation symmetry.
@@
$k_{e}$ specifies an edge region and is chosen as $k_{e}=[L_{q}/4]$ for numerical stability when the energy gap is small or equivalently the correlation length is large. We can numerically confirm $Q_{\rm left}+Q_{\rm right}=0$ holds because of the charge neutrality in the entire system, although $\delta n_{k}$ itself is not inversion symmetric. (Typically, $|Q_{\rm left}+Q_{\rm right}|=O(10^{-8\sim-7})$ for the system size $L=600$.) Note that the present excess edge charges are different from the (smeared or weighted) edge charges widely discussed in the literatures  Resta and Vanderbilt (2007); Vanderbilt (2018); Vanderbilt and King-Smith (1993); Watanabe et al. (2021); Furuya and Sato (2021). The smeared edge charge exactly equals the polarization by definition, while our excess edge charge only partly contributes to the polarization as will be discussed in the following. Our excess charge in the present lattice model would correspond to the extra charge in Bloch electron systems discussed in the previous study Rhim et al. (2017).
@@
where $P_{x}$ is the polarization evaluated with use of the Resta formula under the open boundary condition. Note that $P_{x}$ under the open boundary condition numerically coincides with the one under the periodic boundary condition, which would be a general property as discussed in Appendix C. We confirm that typical difference between $\braket{\hat{P}_{x}},(P_{\rm bulk}+P_{\rm edge})$, and $P_{x}$ is small ($O(10^{-4\sim-3})$ when $L=600$), and it decreases as the system size increases in the numerical calculations. Although the total polarization is a quantized constant under the symmetry, each of $P_{\rm bulk}$ and $P_{\rm edge}$ changes continuously with the parameters. The bulk contribution decreases as the modulation strength is reduced and approaches zero, $P_{\rm bulk}\to 0$, when $\delta\to 0$ or $\Delta\to 0$, because the system becomes a simple metal with a uniform bulk charge density in that limit. At the same time, the edge contribution increases and approaches the quantized value, which means that the excess edge charges $Q_{\rm left/right}$ become quantized in the weak modulation limit under the $I_{k}$-symmetry. Similar quantization was discussed previously, but the fundamental difference is that the edge charges in the previous studies are defined so that they are equal to the total polarization for any parameters Resta and Vanderbilt (2007); Vanderbilt (2018); Vanderbilt and King-Smith (1993); Watanabe et al. (2021); Furuya and Sato (2021). In contrast to the weak modulation limit, the total polarization is dominated by the bulk contribution while the edge contribution is decreased when the modulation is large, because the system is deep in an insulating regime. In the strong modulation limit ($|\delta|\to\infty$ or $|\Delta|\to\infty$), the edge contribution vanishes, $P_{\rm edge}\to 0$. These behaviors are generally seen except for special cases which will be discussed in the next section.

### v2

@@ Abstract
We discuss the charge polarization in a generalized Rice-Mele model at arbitrary particle filling per site as a model of charge ordered systems in one dimension. The model possesses neither the conventional bond-centered inversion symmetry nor the one site translation symmetry alone, but has combinations of these symmetries. We show that the charge polarization in the ground state is quantized by the combined symmetry and is characterized solely by the filling. Especially, the polarization can be $1/2$ (mod 1) in the zero filling limit. Under the open boundary condition, there exist excess charges accumulated near edges of the system irrespective of existence or absence of edge modes. Correspondingly, we decompose the polarization into a bulk contribution and an edge contribution, and numerically demonstrate that the polarization is dominated by the former (latter) when the energy gap is small (large). We also discuss a simple generalization of our model and examine absence/existence of a gapless edge mode protected by the inversion symmetry by introducing intra unit cell and inter unit cell contributions of the charge polarization.
@@
An insulator can have polarization which characterizes asymmetry of a charge configuration. A classical example is the ionic insulator where positive point charges and negative point charges are alternatively distributed. Careful discussions are necessary for an appropriate definition of the polarization even for such a simple system under the periodic boundary condition. It is now widely accepted that the polarization can well be described by the Berry-Zak formula, which clarifies an intimate relationship of the polarization to topology and symmetry  Resta and Vanderbilt (2007); Vanderbilt (2018); Spaldin (2012); King-Smith and Vanderbilt (1993); Vanderbilt and King-Smith (1993); Zak (1989). The celebrated prototypical model of dipole topological insulators is the Su-Schrieffer-Heeger (SSH) model, where the hopping integral has a staggered modulation Su et al. (1979); Su and Schrieffer (1981); János K. Asbóth (2016). The polarization of the SSH model is quantized by an inversion symmetry, and there have been extensive theoretical studies of the generalized SSH models and many interesting properties have been clarified Li et al. (2014); Yao et al. (2017); Pérez-González et al. (2019); Scollon and Kennett (2020); Guo and Chen (2015); Martinez Alvarez and Coutinho-Filho (2019); Huda et al. (2020); Maffei et al. (2018); Xie et al. (2019); Hetényi et al. (2021). Topological nature of the charge polarization is clearly seen in charge pumping where a change of the polarization in an adiabatic cycle in a parameter space is quantized by the Chern number Thouless (1983); Citro and Aidelsburger (2023). Indeed, the Rice-Mele (RM) model Rice and Mele (1982) was experimentally realized in cold atoms and the Thouless pumping was observed Lohse et al. (2016); Nakajima et al. (2016, 2021). In this way, the SSH and RM models are regarded as one of the most important models in the context of topological insulators.
@@
The very basic physics behind the SSH and RM models is charge ordering due to the Peierls instability  Kagoshima (1981); van Smaalen (2004) as they were originally proposed for polyacetylene and diatomic polymers Su et al. (1979); Su and Schrieffer (1981); Rice and Mele (1982). A trimerized variant of the SSH model was also studied with an application to the Peierls system TTF-TCNQ Su and Schrieffer (1981), although related models have been discussed in relation to various systems such as engineered materials and cold atoms Guo and Chen (2015); Martinez Alvarez and Coutinho-Filho (2019); Huda et al. (2020); Maffei et al. (2018); Xie et al. (2019); Hetényi et al. (2021). In a general one-dimensional metal with the Fermi wavenumer $k_{F}$, the charge susceptibility shows singularity at the wavenumber $Q=2k_{F}$ and the system has strong instability toward a charge order. The Peierls phase transition and charge orders have been discussed in both organic materials and inorganic materials, and there are a variety of charge orders with different modulations $Q$  Kagoshima (1981); van Smaalen (2004). In this context, the standard SSH and RM models correspond to the simplest charge order with $Q=\pi$. In spite of the extensive studies on the charge polarization, however, focus has been made mainly on this particular case and polarizations for other charge orders with general $Q$-vectors have not been well explored. Since there are a lot of charge ordered one-dimensional materials Kagoshima (1981); van Smaalen (2004), it is natural to consider generalization of the $Q=\pi$ order to $Q=2k_{F}\neq\pi$ orders. Such generalization may be interesting also in the context of symmetry protected topological phases beyond the non-interacting fermions. For example, an $S=1/2$ trimer spin chain exhibits a 1/3 magnetization plateau implying a uniquely gapped ground state under a magnetic field, which can be regarded as a generalization of more familiar spin dimer orders Hida (1994); Hu et al. (2014); Bera et al. (2022). Modulated spin systems could have “charge polarization” associated with the U(1) spin rotation symmetry similarly to charge ordered fermion systems Oshikawa et al. (1997); Nakamura and Todo (2002); Tasaki (2018); Watanabe et al. (2021); Furuya and Sato (2021).
@@ II Quantized charge polarization under symmetry
where the sites are labeled as $j=0,1,2,\cdots,L-1$. See Fig. 1. The modulation wavenumber is given by $Q=2\pi\nu$, where $\nu=N/L=p/q$ is the particle filling per site with $N$ being the total number of fermions and $p,q$ being coprime integers ($p$ fermions per $q$-site unit cell on average). The system size is assumed to be an integer multiple of the unit cell for both the periodic and open boundary conditions, $L=L_{q}q$, and is taken to be $L=600$ for numerical calculations in this study. The modulation could be induced by interactions or arise from different atom species in real materials, but here $(\delta,\Delta)$ are model parameters of the system. The modulation factor $\cos(Qj)$ reduces to the conventional factor $(-1)^{j}$ at the half-filling $\nu=1/2$. One can introduce a phase shift such as $\cos(Q(j-j_{0}))$ with $j_{0}=1,2,\cdots,q-1$ for which the corresponding Hamiltonian is related to the above $H_{0}$ by the $j_{0}$-site translation (see Sec. IV). It is straightforward to numerically calculate single-particle spectra of Eq. (1) and there exists an energy gap $\Delta E$ between the $N$-th and $(N+1)$-th single-particle energy eigenvalues under the periodic boundary condition as shown in Fig. 2. One can add interactions such as $H_{\rm int}=\sum_{ij}V_{ij}n_{i}n_{j}$ where $n_{j}=c_{j}^{\dagger}c_{j}$ and $V_{ij}$ keeps the symmetry of the system. A large part of our discussions including symmetry-based arguments presented below is applicable also to interacting systems, which will be important when one considers bosonic particles and spin systems. Note that our ground state is a uniquely gapped insulator with an integer filling per unit cell, which is distinct from interaction-driven degenerate ground states Aligia and Ortiz (1999).
Figure 1: Schematic pictures of the system for (a) the periodic boundary condition and (b) the open boundary condition. $I_{0}$ is the inversion about the bond $\braket{L-1,0}$ and $T_{x}$ is the one site translation. The open boundary condition is realized by cutting the hopping integral on the bond $\braket{L-1,0}$ in the periodic boundary condition.
@@
Let us discuss symmetry of the Hamiltonian (1) and the resulting quantization of the charge polarization under the periodic boundary condition. Obviously, neither the one site translation symmetry $T_{x}:c_{j}\to c_{j+1}$ nor the bond-centered inversion symmetry $I_{0}:c_{j}\to c_{L-j-1}$ about the bond $\braket{L-1,0}$ alone is preserved for general $Q$ when $\delta$ and $\Delta$ are non-zero. ($\braket{j,j+1}$ denotes the bond between the sites $j$ and $j+1$.) However, one can see that appropriate combinations of these symmetries are preserved. Indeed, the Hamiltonian is invariant under the combined symmetry when either $\delta$ or $\Delta$ is zero,
@@
The summand $(n_{j}-\nu)$ corresponds to the total electric charge at the site $j$ including the uniform ion contribution $-\nu$. We note that the Resta formula itself is well-defined and can be used for both the periodic boundary condition and the open boundary condition. The operator $U_{x}$ is often called Lieb-Schultz-Mattis twist operator  Oshikawa et al. (1997); Nakamura and Todo (2002); Bohm (1949); Lieb et al. (1961); Tasaki (2018, ) and its expectation value in a uniquely gapped ground state is known to be $|\braket{U_{x}}|\simeq 1$ in one dimension. This is because the energy of the variational state |$\ket{\Phi}=U_{x}\ket{\Psi}$ converges to the ground state energy in the thermodynamic limit and therefore |$\ket{\Phi}$ must be essentially proportional to the ground state with a phase factor, |$\ket{\Phi}\simeq e^{i2\pi P_{x}}\ket{\Psi}$, in a uniquely gapped system. See also Appendix A. Thus the Resta formula for the polarization is well-defined in one dimensional uniquely gapped ground state, although the expectation value is vanishing in the thermodynamic limit for general dimensions and an appropriate modification is necessary to resolve such a problem  Watanabe and Oshikawa (2018); Tada and Oshikawa (2023).
@@
Note that these properties themselves are true for both the periodic boundary condition and the open boundary condition, since they are just algebraic relations of the operators and independent of a quantum state within the fixed filling Hilbert space. The above equations imply that, for the uniquely gapped ground state |$\ket{\Psi}$ which is an eigenstate of the combined operator $I_{k}$ under the periodic boundary condition,
@@
In this section, we consider the model Eq. (1) with the sites $j=0,1,\cdots,L-1$ under the open boundary condition where there is no hopping between the right edge site $j=L-1$ and the left edge site $j=0$ (Fig. 1). The inversion acts as $I_{0}:c_{j}\to c_{L-j-1}$, as in the periodic boundary condition. In the standard SSH model ($\delta\neq 0,\Delta=0$) at $\nu=1/2$, there can be gapless edge modes thanks to the chiral symmetry, but existence of such single-particle edge modes are not generally guaranteed in a system only with point group symmetry Song et al. (2017); Huang et al. (2017). Similarly, there are no edge modes in the present system at almost all filling except for some special cases which will be discussed in Sec. IV. Nevertheless, one may naively expect that there are extra charges accumulated near edges of the system for a given filling. Here we introduce excess charges over one unit cell (length $q$) compared to the uniform background charges,
@@
The excess charge $\delta n_{k}$ corresponds to the total charge of electrons and ions contained in the $k$-th unit cell. Figure 6 shows examples of the fermion density profiles $\braket{n_{j}}$ and the corresponding excess charges $\delta n_{k}$. Clearly, the excess charges vanish in the bulk region of the system, although the fermion density itself oscillates with the period $q$. This is a general property for a gapped periodic system which has unit cell translation symmetry.
@@
$k_{e}$ specifies an edge region of length $O(1)$ and is supposed to be sufficiently larger than the correlation length of the system. In the numerical calculations, however, it is chosen as $k_{e}=[L_{q}/4]$ for numerical stability when the energy gap is small or equivalently the correlation length is large. We can numerically confirm $Q_{\rm left}+Q_{\rm right}=0$ holds because of the charge neutrality in the entire system, although $\delta n_{k}$ itself is not inversion symmetric in general. (Typically, $|Q_{\rm left}+Q_{\rm right}|=O(10^{-8\sim-7})$ for the system size $L=600$. The equation $Q_{\rm left}+Q_{\rm right}=0$ is satisfied in the thermodynamic limit, $\lim_{k_{e}\to\infty}\lim_{L\to\infty}(Q_{\rm left}+Q_{\rm right})=0$, because $Q_{k}\simeq 0$ for the $k$-th unit cell in the bulk region in a gapped system.) We emphasize that the present excess edge charges are different from the (smeared or weighted) edge charges widely discussed in the literatures  Resta and Vanderbilt (2007); Vanderbilt (2018); Vanderbilt and King-Smith (1993); Watanabe et al. (2021); Furuya and Sato (2021). The smeared edge charge is given by a convolution of the charge density with a broadening function and it exactly equals the polarization. On the other hand, our excess edge charge only partly contributes to the polarization as will be discussed in the following. Our excess charge in the present lattice model would correspond to the extra charge in Bloch electron systems discussed in the previous study Rhim et al. (2017). It may be regarded as a many-body variant of the previously introduced extra edge charge based on the trivial decomposition of the Resta formula.
@@
where $P_{x}$ is the polarization evaluated with use of the Resta formula under the open boundary condition. (As mentioned in the previous section, the Resta formula is well-defined both for the periodic boundary condition and the open boundary condition.) Note that $P_{x}$ under the open boundary condition numerically coincides with the one under the periodic boundary condition, which would be a general property as discussed in Appendix C. We confirm that typical difference between $\braket{\hat{P}_{x}},(P_{\rm bulk}+P_{\rm edge})$, and $P_{x}$ is small ($O(10^{-4\sim-3})$ when $L=600$), and it decreases as the system size increases in the numerical calculations. Although the total polarization is a quantized constant under the symmetry, each of $P_{\rm bulk}$ and $P_{\rm edge}$ changes continuously with the parameters. The bulk contribution decreases as the modulation strength is reduced and approaches zero, $P_{\rm bulk}\to 0$, when $\delta\to 0$ or $\Delta\to 0$, because the system becomes a simple metal with a uniform bulk charge density in that limit. At the same time, the edge contribution increases and approaches the quantized value, which means that the excess edge charges $Q_{\rm left/right}$ become quantized in the weak modulation limit under the $I_{k}$-symmetry. Similar quantization was discussed previously, but the fundamental difference is that the edge charges in the previous studies are defined so that they are equal to the total polarization for any parameters Resta and Vanderbilt (2007); Vanderbilt (2018); Vanderbilt and King-Smith (1993); Watanabe et al. (2021); Furuya and Sato (2021). In contrast to the weak modulation limit, the total polarization is dominated by the bulk contribution while the edge contribution is decreased when the modulation is large, because the system is deep in an insulating regime. In the strong modulation limit ($|\delta|\to\infty$ or $|\Delta|\to\infty$), the edge contribution vanishes, $P_{\rm edge}\to 0$. These behaviors are generally seen except for special cases which will be discussed in the next section.